OpenStudy (anonymous):

Write an equation for a line perpendicular to y = -3 x + 3 and passing through the point (-6,-4) how do I do this?

OpenStudy (anonymous):

to make a line perpendicular, you take the current slope and times it by -1 and take the recipricle of it. so if my slope was 3, then I do -1*3 = -3, then the recipricle is -1/3 (a fraction) make sense so far?

OpenStudy (anonymous):

Yes because m1*m2 = -1 correct?

OpenStudy (anonymous):

great, so your slope is going to be 1/3, the next part is...

OpenStudy (anonymous):

you mean -1/3?

OpenStudy (anonymous):

to take the standard slope intercept equation form, and plug in our new values: y =mx+b <-slope intercept form we have m, and are given x and y as the 'point the new line must pass through' so substitue in the values that you can right now and show me what you get

OpenStudy (anonymous):

therefore b=6?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

and to answer yout question, it is 1/3. see, we have -3 as m, then we times it by -1, so (-1)*-3 = 3 then we take the recipricle, so 3 becomes a fraction 1/3

OpenStudy (anonymous):

it would be -6 if our slope was -1/3, but our slope needs to be 1/3 inorder to have the new line be perpendicular ya? make sense? ask all you want, i'm here to help you

OpenStudy (anonymous):

Okay so if m1 is -3 then m2 is unknown correct?

OpenStudy (anonymous):

in order to find m2 we multiply -1*-3?

OpenStudy (anonymous):

or do we put it into a fraction?

OpenStudy (anonymous):

m2 means the slope of the new line we are looking for, which is perpendicular?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

"in order to find m2 we multiply -1*-3?" this idea is right, yes

OpenStudy (anonymous):

okay which would in turn = 3 not 1/3

OpenStudy (anonymous):

m2 = m1 * (-1) <- part 1 of finding the slope of a line perpendicular part 2 of finding the slope of a line perpendicular is taking the 'opposite form' of m2 which we got from the part just above. so if part 1 had m2 = 1/4, then part 2 is to take the 'opposite form' or flip m2, like this: 1/4 becomes 4

OpenStudy (anonymous):

here are some examples: m1 = 5 m2 = (-1)*5 = -5 Part 1 then we flip it in part 2 so now m2 = -1/5 another example m1 = -1/8 so, m2 = (m1) * (-1) = (-1/8) * (-1) = 1/8 then flip it 1/8 becomes 8 so m2 = 8

OpenStudy (anonymous):

Oh okay, I see what your saying, that makes sense now. So in turn since we got -3 it would come out too -3/1 which would give us 1/3

OpenStudy (anonymous):

bingo! that is exactly what you do every time you want find the slope of a line perpendicular to a given line

OpenStudy (anonymous):

so my equation would look like y=1/3x +2 ?

OpenStudy (anonymous):

yes, all except I got -2 not 2, let me double check...

OpenStudy (anonymous):

yes it is -2 I messed up on my math

OpenStudy (anonymous):

so, when solving for b, your equation should look like this: -4 = (1/3) * (-6) + b

OpenStudy (anonymous):

oh ok ok ya you got it!

OpenStudy (anonymous):

so just to re-cap, when ever you want to find the slope of a perpendicular line, you times it by -1 and flip it.

OpenStudy (anonymous):

okay and for a parallel line its m1=m2?

OpenStudy (anonymous):

yes, you got it!

OpenStudy (anonymous):

I wrote m1*m2=-1 for a perpendicular line.

OpenStudy (anonymous):

That's what the professor wrote down.

OpenStudy (anonymous):

ya, he is right

OpenStudy (anonymous):

then you just have to flip m2

OpenStudy (anonymous):

Could you show an example so if we have m1 = -3 and m2= 1/3 what would you do?

OpenStudy (anonymous):

oh never-mind just answered my own question!

OpenStudy (anonymous):

Well you have been great help thank you very much!!!

OpenStudy (anonymous):

you are welcome! good job on asking questions

OpenStudy (anonymous):

Yeah thanks, I love this site such a big help for me sometimes. Well have a good night appreciate the help.